Find the value of Charge q0 if its electric field has a value of 8.45 x 10-3 N/C and a force F0 of 1.3 N


1
Expert's answer
2022-05-20T04:53:02-0400
"F=qE"

"q=\frac{F}{E}=\frac{1.3}{8.45*10^{-3}}=154\:\rm C"